# Enabling Gigantic MIMO Beamforming

### v1

@@ II Microwave Linear Analog Computer
III Analog Computing for
Regularized Least Squares
In [15], we have shown how a MiLAC can be used to compute the linear minimum mean square error (LMMSE) estimator of a linear observation process with reduced computational complexity. In this section, we demonstrate that a MiLAC can also efficiently compute RLS in the analog domain, an operation of particular interest in wireless communications. This is achieved by first recalling the concept of LMMSE estimator and showing that it includes RLS as a special case.
Consider a linear observation process $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$, where $\mathbf{y}\in\mathbb{C}^{Y\times 1}$ is the observation vector, $\mathbf{H}\in\mathbb{C}^{Y\times X}$ is a known constant matrix, $\mathbf{x}\in\mathbb{C}^{X\times 1}$ is the unknown random vector with mean $\bar{\mathbf{x}}=\mathbf{0}_{X\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{x}}$, and $\mathbf{n}\in\mathbb{C}^{Y\times 1}$ is the random noise vector with mean $\bar{\mathbf{n}}=\mathbf{0}_{Y\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{n}}$. As discussed in [15], the LMMSE estimator of $\mathbf{x}$ is equivalently given by
when $\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}+\mathbf{C}_{\mathbf{x}}^{-1}$ and $\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}$ are both invertible, which holds true in general.
This LMMSE estimator has been obtained with no assumptions on the covariance matrices $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ beyond their invertibility. With specific assumptions on $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$, we can show that it includes the RLS, also known as ridge regression, as a special case of practical interest. Specifically, assume that the entries of $\mathbf{x}$ and $\mathbf{n}$ are uncorrelated and have equal variance, i.e., $\mathbf{C}_{\mathbf{x}}=\lambda_{\mathbf{x}}\mathbf{I}_{X}$ and $\mathbf{C}_{\mathbf{n}}=\lambda_{\mathbf{n}}\mathbf{I}_{Y}$, with $\lambda_{\mathbf{x}},\lambda_{\mathbf{n}}\in\mathbb{R}$ and $\lambda_{\mathbf{x}},\lambda_{\mathbf{n}}>0$. Thus, by introducing $\lambda\in\mathbb{R}$ such that $\lambda=\lambda_{\mathbf{n}}/\lambda_{\mathbf{x}}$, (8) and (9) simplify as
respectively, which are known as the expressions of RLS. Interestingly, the expressions in (10) and (11) are widely used in MIMO wireless communications to realize the so-called R-ZFBF at the transmitter and MMSE detection at the receiver, as effective linear transmission and reception techniques, respectively [16, Chapters 7, 12].
@@ III Analog Computing for
The expressions of RLS can be efficiently returned on the output vector $\mathbf{v}_{2}$ of a MiLAC having $N=Y$ input ports and $M=X$ output ports with no input, i.e., $P=X+Y$ total ports. This is achieved by exploiting (4) and (6), as discussed for the LMMSE estimator in [15]. The input vector of the MiLAC is set as $\mathbf{u}=\mathbf{y}$ and the tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$ are set according to (7) such that the blocks of the matrix $\mathbf{P}$ give the desired expression in $\mathbf{v}_{2}$. In detail, the RLS expressions in (10) and (11) can be computed by setting the matrix $\mathbf{P}$ as illustrated in Tabs. I and II.
TABLE II: MiLAC output $\mathbf{v}_{2}$ for different $\mathbf{P}$ with $\mathbf{P}_{22}$ invertible.
@@
Remark 1.
RLS, or ridge regression, can be efficiently computed using a MiLAC, which offers significantly reduced computational complexity compared to a digital computer. This is because a MiLAC directly performs matrix-matrix multiplications and matrix inversion operations in the analog domain. On the one hand, the computational complexity of computing RLS with a MiLAC is driven by the complexity of computing (7) for a given matrix $\mathbf{P}$. Thus, the complexity required to compute RLS with a MiLAC is the same as the complexity of computing the LMMSE estimator, i.e., $6XY$ real operations, as discussed in [15]. On the other hand, a digital computer calculates the RLS with a complexity of $8\min\{X^{2}Y+X^{3}/3,XY^{2}+Y^{3}/3\}$ real operations as it requires one matrix-matrix product and a matrix inversion (see [15, Appendix] for more details on the complexity of a matrix-matrix product and a matrix inversion).
@@ IV A Novel Beamforming Strategy:
In this section, we show how a MiLAC can be used to implement a novel beamforming technique, that we denote as MiLAC-aided beamforming, at the transmitter as well as receiver side. MiLAC-aided beamforming offers circuit and computational complexity benefits by fully operating in the analog domain. To illustrate MiLAC-aided beamforming, we consider a MIMO communication system between $N_{T}$ transmitting antennas and $N_{R}$ receiving antennas through the wireless channel $\mathbf{H}\in\mathbb{C}^{N_{R}\times N_{T}}$. We denote as $\mathbf{x}\in\mathbb{C}^{N_{T}\times 1}$ the transmitted signal and as $\mathbf{y}\in\mathbb{C}^{N_{R}\times 1}$ the received signal such that $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$, where $\mathbf{n}\in\mathbb{C}^{N_{R}\times 1}$ is the additive white Gaussian noise (AWGN)  [^1] We consider here a general MIMO system, such as single-user MIMO or multi-user with one symbol per user. Both single- and multi-user systems will be specifically studied in the following subsections..
@@
Figure 2: MiLAC-aided beamforming at the transmitter, performing R-ZFBF in the analog domain.
IV-A R-ZFBF at the Transmitter
@@
Assume that the MIMO system $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$ is a multi-user system between an $N_{T}$-antenna transmitter and $N_{R}$ single-antenna users, with $N_{R}\leq N_{T}$. The signal received at the $N_{R}$ users $\mathbf{y}$ is used to detect the symbols $\mathbf{s}\in\mathbb{C}^{N_{S}\times 1}$, with $N_{S}=N_{R}$, and the transmitted signal is given by $\mathbf{x}=\mathbf{W}\mathbf{s}$, where $\mathbf{W}\in\mathbb{C}^{N_{T}\times N_{R}}$ is the precoding matrix. Such a transmitter can be implemented purely in the analog domain by a MiLAC having $N=N_{S}$ ports receiving the symbol vector in input and $M=N_{T}$ ports delivering the output to antennas perfectly matched to the reference impedance $Z_{0}=Y_{0}^{-1}$, as represented in Fig. 2. Since the antennas are perfectly matched, their Thevenin equivalent circuit is an impedance $Z_{0}=Y_{0}^{-1}$ connected to ground [17, Chapter 2.13]. Thus, the MiLAC in Fig. 2 behaves as when its last $N_{T}$ ports are connected to ground through an impedance $Z_{0}=Y_{0}^{-1}$, i.e., as analyzed in Section II. By exploiting Tab. II, it is possible to reconfigure the microwave network of the MiLAC to efficiently execute R-ZFBF [16, Chapter 12]. Specifically, a MiLAC computing RLS can realize R-ZFBF, with precoding matrix $\mathbf{W}_{\text{R-ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H}+\lambda\mathbf{I}_{N_{R}})^{-1}$, where $\lambda$ is an arbitrary regularization factor. This can be obtained by reconfiguring the tunable admittance components of the MiLAC according to the channel $\mathbf{H}$ on a per coherence block basis, and do not require any operation on a per-symbol basis.
@@
According to Remark 1, R-ZFBF can be achieved with only $6N_{T}N_{R}$ real operations, which are necessary to set the tunable admittance components of the MiLAC. Conversely, the computational complexity of digitally calculating the R-ZFBF matrix is $8(N_{T}N_{R}^{2}+N_{R}^{3}/3)$ real operations. This confirms the significant benefits of MiLAC-aided beamforming in terms of computational complexity.
Figure 3: MiLAC-aided beamforming at the receiver, performing MMSE in the analog domain.
IV-B MMSE at the Receiver
As discussed on the transmitter side, MiLAC-aided beamforming can also be introduced on the receiver side. Assume that the MIMO system $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$ is a single-user system between an $N_{T}$-antenna transmitter and a $N_{R}$-antenna receiver, with $N_{R}\geq N_{T}$. The transmitted signal $\mathbf{x}$ contains the symbols to be detected at the receiver, i.e., $\mathbf{x}=\mathbf{s}$, and the receiver uses the signal $\mathbf{z}=\mathbf{G}\mathbf{y}\in\mathbb{C}^{N_{T}\times 1}$ to accurately detect the symbols, where $\mathbf{G}\in\mathbb{C}^{N_{T}\times N_{R}}$ is the combining matrix. This receiver can be implemented with MiLAC having $N=N_{R}$ ports acquiring the input from the receiving antennas and $M=N_{S}$ ports providing the output vector $\mathbf{z}$, as represented in Fig. 3. Note that in Fig. 3 we assume the antennas to be perfectly matched to the reference impedance $Z_{0}=Y_{0}^{-1}$, such that their Thevenin equivalent circuit is a voltage generator (inducing a voltage $y_{n}$ at port $n$, for $n=1,\ldots,N$) in series to an impedance $Z_{0}=Y_{0}^{-1}$ [17, Chapter 2.13], i.e., as in the MiLAC analyzed in Section II. By using Tab. I, it is possible to reconfigure the microwave network of the MiLAC to efficiently the MMSE receiver [16, Chapter 7]. Specifically, a MiLAC computing RLS can realize the MMSE receiver with combining matrix $\mathbf{G}_{\text{MMSE}}=(\mathbf{H}^{H}\mathbf{H}+\lambda\mathbf{I}_{N_{T}})^{-1}\mathbf{H}^{H}$. If the covariance matrix of the transmitted signal $\mathbf{x}$ is $\mathbf{C}_{\mathbf{x}}=P_{T}/N_{T}\mathbf{I}_{N_{T}}$, where $P_{T}$ is the transmit power, and the covariance matrix of the noise $\mathbf{n}$ is $\mathbf{C}_{\mathbf{n}}=\sigma_{\mathbf{n}}^{2}\mathbf{I}_{N_{R}}$, where $\sigma_{\mathbf{n}}^{2}$ is the noise power, the optimal regularization factor $\lambda$ is given by $\lambda=N_{T}\sigma_{\mathbf{n}}^{2}/P_{T}$ [16, Chapter 7]. This can be realized by solely reconfiguring the MiLAC on a per coherence block basis, while no operation is required on a per-symbol basis.

### v2

@@ III-B5 Ordinary Matched Filtering
GLS $\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mathbf{0}_{X}\end{bmatrix}$
GMF $\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
@@
OLS $\left(\mathbf{H}^{H}\mathbf{H}\right)^{-1}\mathbf{H}^{H}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{I}_{Y}&\mathbf{H}\\ \mathbf{H}^{H}&\mathbf{0}_{X}\end{bmatrix}$
OMF $\lambda^{-1}\mathbf{H}^{H}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{I}_{Y}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\lambda\mathbf{I}_{X}\end{bmatrix}$
III-C Computing the Special Cases of the LMMSE Estimator with a MiLAC
@@
The expressions of the five special cases of the LMMSE estimator can be efficiently returned on the output vector $\mathbf{v}_{2}$ of a MiLAC having $N=Y$ input ports and $M=X$ output ports with no input, i.e., $P=X+Y$ total ports. This is achieved by exploiting (4) and (6), as discussed for the LMMSE estimator in Part I of this paper. The input vector of the MiLAC is set as $\mathbf{u}=\mathbf{y}$ and the tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$ are set according to (7) such that the blocks of the matrix $\mathbf{P}$ give the desired expression in $\mathbf{v}_{2}$. In detail, the expressions of the special cases in (10)-(17) can be computed by setting the matrix $\mathbf{P}$ as illustrated in Tabs. II and III, depending on the invertibility of the blocks $\mathbf{P}_{11}$ and $\mathbf{P}_{22}$. When the block $\mathbf{P}_{11}$ is invertible, $\mathbf{v}_{2}$ is given by (4), as reported in Tab. II, while when the block $\mathbf{P}_{22}$ is invertible, $\mathbf{v}_{2}$ is given by (6), as reported in Tab. III. Note that GLS and OLS are computed via a matrix $\mathbf{P}$ whose blocks $\mathbf{P}_{11}$ and $\mathbf{P}_{22}$ are not both invertible. Thus, for GLS and OLS the expressions in Tabs. II and III are not equivalent in general, while they are equivalent for the LMMSE estimator and the other special cases.
Remark 1.
The five special cases of the LMMSE estimator can be efficiently computed using a MiLAC, which offers significantly reduced computational complexity compared to a digital computer. This is because a MiLAC directly performs matrix-matrix multiplications and matrix inversion operations in the analog domain. On the one hand, the computational complexity of computing the special cases with a MiLAC is driven by the complexity of computing (7) for a given matrix $\mathbf{P}$. Thus, the complexity required to compute GLS, RLS, and OLS with a MiLAC is the same as the complexity of computing the LMMSE estimator, i.e., $6XY$ real operations, as discussed in Part I of this paper. Besides, computing (7) for the GMF and OMF require approximately $2XY$ real operations for both cases $i\neq k$ and $i=k$, resulting in a total of $4XY$ real operations, where we assumed that the operations involving $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ are precomputed offline. On the other hand, a digital computer calculates the GLS with the same complexity as the LMMSE estimator due to the two matrix-matrix products and matrix inversion, i.e., $8(XY^{2}+X^{2}Y+\min\{X^{3},Y^{3}\}/3)$, as discussed in Part I of this paper. The RLS and OLS can be computed with reduced complexity as they require only one matrix-matrix product and a matrix inversion, namely $8\min\{X^{2}Y+X^{3}/3,XY^{2}+Y^{3}/3\}$. Besides, the GMF and OMF only require three and one matrix-vector products, respectively, resulting in a complexity of $8(X^{2}+XY+Y^{2})$ and $8XY$ real operations, respectively.
TABLE III: MiLAC output $\mathbf{v}_{2}$ for different $\mathbf{P}$ with $\mathbf{P}_{22}$ invertible.
@@
GLS $\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\right)^{-1}\mathbf{u}$ $\begin{bmatrix}\mathbf{0}_{Y}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
GMF $\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
@@
OLS $\mathbf{H}^{H}\left(\mathbf{H}\mathbf{H}^{H}\right)^{-1}\mathbf{u}$ $\begin{bmatrix}\mathbf{0}_{Y}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{I}_{X}\end{bmatrix}$
OMF $\lambda^{-1}\mathbf{H}^{H}\mathbf{u}$ [±$\begin{bmatrix}\pm\lambda\mathbf{I}_{Y}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\mathbf{I}_{X}\end{bmatrix}$
@@ IV A Novel Beamforming Strategy:
In this section, we show how a MiLAC can be used to implement a novel beamforming technique, that we denote as MiLAC-aided beamforming, at the transmitter as well as receiver side. MiLAC-aided beamforming has the same flexibility as digital beamforming, while it offers circuit and computational complexity benefits by fully operating in the analog domain. To illustrate MiLAC-aided beamforming, we consider a MIMO communication system between $N_{T}$ transmitting antennas and $N_{R}$ receiving antennas through the wireless channel $\mathbf{H}\in\mathbb{C}^{N_{R}\times N_{T}}$. We denote as $\mathbf{x}\in\mathbb{C}^{N_{T}\times 1}$ the transmitted signal and as $\mathbf{y}\in\mathbb{C}^{N_{R}\times 1}$ the received signal such that $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$, where $\mathbf{n}\in\mathbb{C}^{N_{R}\times 1}$ is the additive white Gaussian noise (AWGN)  [^1] We consider here a general MIMO system, such as single-user MIMO or multi-user with one symbol per user. Both single- and multi-user systems will be specifically studied in the following subsections.. We begin by showing how MiLAC-aided beamforming can be used to implement an arbitrary beamforming matrix.
@@
Figure 2: MiLAC-aided beamforming at the transmitter, performing $\mathbf{x}=\mathbf{W}\mathbf{s}$ in the analog domain, where $\mathbf{s}$ is the symbol vector and $\mathbf{W}$ is an arbitrary precoding matrix.
IV-A Arbitrary Beamforming
@@
Consider a transmitter that linearly precodes a vector of $N_{S}$ symbols $\mathbf{s}\in\mathbb{C}^{N_{S}\times 1}$ to obtain the transmitted signal as $\mathbf{x}=\mathbf{W}\mathbf{s}$, where $\mathbf{W}\in\mathbb{C}^{N_{T}\times N_{S}}$ is an arbitrary precoding matrix. Such a transmitter can be implemented purely in the analog domain by a MiLAC having $N=N_{S}$ ports receiving the symbol vector in input and $M=N_{T}$ ports delivering the output to antennas perfectly matched to the reference impedance $Z_{0}=Y_{0}^{-1}$, as represented in Fig. 2. Since the antennas are perfectly matched, their Thevenin equivalent circuit is an impedance $Z_{0}=Y_{0}^{-1}$ connected to ground [36, Chapter 2.13]. Thus, the MiLAC in Fig. 2 behaves as when its last $N_{T}$ ports are connected to ground through an impedance $Z_{0}=Y_{0}^{-1}$, i.e., as analyzed in Section II. Following the analysis in Section II, such a MiLAC returns $\mathbf{x}=\mathbf{W}\mathbf{s}$ on its last $N_{T}$ ports when the signal $\mathbf{s}$ is fed as input on the first $N_{S}$ ports and the tunable admittance components are set such that the matrix $\mathbf{P}$ is
$\mathbf{P}=\begin{bmatrix}\mathbf{I}_{N_{S}}&\mathbf{0}_{N_{S}\times N_{T}}\\ -\mathbf{W}&\mathbf{I}_{N_{T}}\end{bmatrix},$ (18)
according to (4) and (6). To this end, the tunable admittance components are set based on (7) where $\mathbf{P}$ is given by (18), yielding
$Y_{i,k}=\begin{cases}Y_{0}\left[\mathbf{W}\right]_{i-N,k}&N_{S}<i,\;k\leq N_{S}\\ 0&\text{Otherwise}\end{cases},$ (19)
for $i\neq k$, and
$Y_{k,k}=\begin{cases}-Y_{0}\sum_{n_{T}=1}^{N_{T}}\left[\mathbf{W}\right]_{n_{T},k}&k\leq N_{S}\\ 0&k>N_{S}\end{cases},$ (20)
for $k=1,\ldots,N_{S}+N_{T}$. From (19) and (20), we observe that the tunable admittance components of a MiLAC can be reconfigured in closed form to implement any arbitrary beamforming $\mathbf{W}$ matrix in the analog domain. The computational complexity of this beamforming technique, denoted as MiLAC-aided beamforming, is given by the complexity of designing $\mathbf{W}$ (depending on the adopted precoding strategy) and computing (19) and (20) (requiring in total $4N_{S}N_{T}$ real operations). Note that these operations are executed at each coherence block, while no operation is required on a per-symbol basis since the matrix-vector product $\mathbf{W}\mathbf{s}$ is performed in the analog domain.
Figure 3: MiLAC-aided beamforming at the receiver, performing $\mathbf{z}=\mathbf{G}\mathbf{y}$ in the analog domain, where $\mathbf{y}$ is the received signal and $\mathbf{G}$ is an arbitrary combining matrix.
As discussed on the transmitter side, MiLAC-aided beamforming can also be introduced on the receiver side. Specifically, a MiLAC can be used to linearly combine the received signal $\mathbf{y}\in\mathbb{C}^{N_{R}\times 1}$ to obtain the signal used for detection $\mathbf{z}\in\mathbb{C}^{N_{S}\times 1}$ as $\mathbf{z}=\mathbf{G}\mathbf{y}$, where $\mathbf{G}\in\mathbb{C}^{N_{S}\times N_{R}}$ is an arbitrary combining matrix. This can be obtained with MiLAC having $N=N_{R}$ ports acquiring the input from the receiving antennas and $M=N_{S}$ ports providing the output vector $\mathbf{z}$, as represented in Fig. 3. Note that in Fig. 3 we assume the antennas to be perfectly matched to the reference impedance $Z_{0}=Y_{0}^{-1}$, such that their Thevenin equivalent circuit is a voltage generator (inducing a voltage $y_{n}$ at port $n$, for $n=1,\ldots,N$) in series to an impedance $Z_{0}=Y_{0}^{-1}$ [36, Chapter 2.13], i.e., as in the MiLAC analyzed in Section II.
Among the possible linear precoding and combining strategies, MiLAC-aided beamforming can efficiently perform R-ZFBF, ZFBF, and MBF at the transmitter and MMSE, ZF, and MF at the receiver by exploiting the capability of a MiLAC to compute the corresponding special case of the LMMSE estimator. This leads to remarkable computational complexity benefits, as illustrated in the following.
IV-B LMMSE-Inspired Beamforming at the Transmitter:
R-ZFBF, ZFBF, and MBF Transmitters
Assume that the previously considered MIMO system $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$ is a multi-user system between an $N_{T}$-antenna transmitter and $N_{R}$ single-antenna users, with $N_{R}\leq N_{T}$. The signal received at the $N_{R}$ users $\mathbf{y}$ is used to detect the symbols $\mathbf{s}\in\mathbb{C}^{N_{S}\times 1}$, with $N_{S}=N_{R}$, and the transmitted signal is given by $\mathbf{x}=\mathbf{W}\mathbf{s}$, where $\mathbf{W}\in\mathbb{C}^{N_{T}\times N_{R}}$ is the precoding matrix implemented by a MiLAC, as in Fig. 2. By exploiting the operations included in Tab. III, it is possible to reconfigure the microwave network of the MiLAC to efficiently execute three popular linear transmitters [35, Chapter 12]. First, a MiLAC computing RLS can realize R-ZFBF, with precoding matrix $\mathbf{W}_{\text{R-ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H}+\lambda\mathbf{I}_{N_{R}})^{-1}$, where $\lambda$ is an arbitrary regularization factor. Second, ZFBF can be implemented by a MiLAC computing OLS, whose precoding matrix is $\mathbf{W}_{\text{ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H})^{-1}$. Third, a MiLAC computing OMF can realize MBF, with precoding matrix given by $\mathbf{W}_{\text{MBF}}=\lambda^{-1}\mathbf{H}^{H}$. These three transmitters are obtained by reconfiguring the tunable admittance components of the MiLAC according to the channel $\mathbf{H}$ on a per coherence block basis, and do not require any operation on a per-symbol basis.
@@ IV-B LMMSE-Inspired Beamforming at the Transmitter:
According to Remark 1, the R-ZFBF and ZFBF transmitters can be achieved with only $6N_{T}N_{R}$ real operations, while the MBF transmitter with only $4N_{T}N_{R}$ real operations, which are necessary to set the tunable admittance components of the MiLAC. Conversely, the computational complexity of digitally calculating the R-ZFBF and ZFBF transmitters is $8(N_{T}N_{R}^{2}+N_{R}^{3}/3)$ real operations, while for the MBF transmitter is $8N_{T}N_{R}$. This confirms the significant benefits of MiLAC-aided beamforming in terms of computational complexity, especially in implementing the R-ZFBF and ZFBF transmitters.
IV-C LMMSE-Inspired Beamforming at the Receiver:
MMSE, ZF, and MF Receivers
Assume that the MIMO system $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$ is a single-user system between an $N_{T}$-antenna transmitter and a $N_{R}$-antenna receiver, with $N_{R}\geq N_{T}$. The transmitted signal $\mathbf{x}$ contains the symbols to be detected at the receiver, i.e., $\mathbf{x}=\mathbf{s}$, and the receiver uses the signal $\mathbf{z}=\mathbf{G}\mathbf{y}\in\mathbb{C}^{N_{T}\times 1}$ to accurately detect the symbols, where $\mathbf{G}\in\mathbb{C}^{N_{T}\times N_{R}}$ is the combining matrix imposed by the MiLAC, as in Fig. 3. By using the operations in Tab. II, it is possible to reconfigure the microwave network of the MiLAC to efficiently implement three widely used receivers [35, Chapter 7]. First, a MiLAC computing RLS can realize the MMSE receiver with combining matrix $\mathbf{G}_{\text{MMSE}}=(\mathbf{H}^{H}\mathbf{H}+\lambda\mathbf{I}_{N_{T}})^{-1}\mathbf{H}^{H}$. If the covariance matrix of the transmitted signal $\mathbf{x}$ is $\mathbf{C}_{\mathbf{x}}=P_{T}/N_{T}\mathbf{I}_{N_{T}}$, where $P_{T}$ is the transmit power, and the covariance matrix of the noise $\mathbf{n}$ is $\mathbf{C}_{\mathbf{n}}=\sigma_{\mathbf{n}}^{2}\mathbf{I}_{N_{R}}$, where $\sigma_{\mathbf{n}}^{2}$ is the noise power, the optimal regularization factor $\lambda$ is given by $\lambda=N_{T}\sigma_{\mathbf{n}}^{2}/P_{T}$ [35, Chapter 7]. Second, the ZF receiver can be realized by a MiLAC computing OLS, whose combining matrix is $\mathbf{G}_{\text{ZF}}=(\mathbf{H}^{H}\mathbf{H})^{-1}\mathbf{H}^{H}$. Third, a MiLAC computing OMF can implement the MF receiver, with combining matrix $\mathbf{G}_{\text{MF}}=\lambda^{-1}\mathbf{H}^{H}$, where the optimal $\lambda$ is $\lambda=N_{T}\sigma_{\mathbf{n}}^{2}/P_{T}$ if $\mathbf{C}_{\mathbf{x}}=P_{T}/N_{T}\mathbf{I}_{N_{T}}$ and $\mathbf{C}_{\mathbf{n}}=\sigma_{\mathbf{n}}^{2}\mathbf{I}_{N_{R}}$ [35, Chapter 7]. These three receivers are realized by solely reconfiguring the MiLAC on a per coherence block basis, while no operation is required on a per-symbol basis.